OpenStudy (anonymous):

DONTTTT UNDERSTAND: write an equation of the line that passes through each pair of points. Does this mean use the formula: y1-2 over x1-x2?

OpenStudy (anonymous):

 iss

OpenStudy (texaschic101):

yes...you would use : slope(m) = (y2 - y1) / (x2 - x1) to find the slope.

OpenStudy (texaschic101):

however, once you find the slope, you will use y - y1 = m(x - x1) to get your equation.

OpenStudy (texaschic101):

do you need an example ?